ROTATIONALLY RESOLVED HIGH-RESOLUTION LASER SPECTROSCOPY OF THE S_{1} ← S_{0} TRANSI-TION OF NAPHTHALENE AND Cl-NAPHTHALENE

\underline{SHUNJI KASAHARA}, RYO YAMAMOTO, KOHEI TADA, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan.

Rotationally resolved high-resolution fluorescence excitation spectra and the Zeeman effects of 0-0 band of S_{1} ← S_{0} electronic transition have been observed for naphthalene, 1-Cl naphthalene (1-ClN), and 2-Cl naphthalene (2-ClN). Sub-Doppler excitation spectra were measured by crossing a single-mode UV laser beam perpendicular to a collimated molecular beam. The typical linewidth was 25 MHz and the absolute wavenumber was calibrated with accuracy 0.0002 cm^{−1} by measurement of the Doppler-free saturation spectrum of iodine molecule and fringe pattern of the stabilized etalon. For naphthalene ^{a} and 2-ClN, the rotationally resolved spectra were obtained, and these molecular constants were determined in high accuracy. The obtained molecular constants of 2-ClN are good agreement with the ones reported by Plusquellic et. al. ^{b} For 1-ClN, the rotational lines were not completely resolved because the fluorescence lifetime is shorter than the one of 2-ClN. Additionally, we have observed the change of the spectra with magnetic field. The Zeeman broadening was mainly observed for the levels of low K_{a} and increasing in proportion to J for given K for both of naphthalene and 2-ClN. The order of magnitude and the J, K-dependence of the observed Zeeman broadening were similar to the other vibronic bands of naphthalene. ^{c}

aD. L. Joo, R. Takahashi, J. O’Reilly, H. Katô, and M. Baba, J. Mol. Spectrosc., 215, 155 (2002).

bD. F. Plusquellic, S. R. Davis, and F. Jahanmir, J. Chem. Phys., 115, 225 (2001).

cH. Katô, S. Kasahara, and M. Baba, Bull. Chem. Soc. Jpn., 80, 456 (2007).